If the amount of CO2 increases in C2H2 + O2 ⇆ CO2 + H2O ΔH = +


1
Expert's answer
2021-06-24T06:19:02-0400

q = m × Cg × ΔT

Molar Mass of CO2 = 44.01

Specific heat capacity of CO2 = 29.86 kJ

∆T= -79 + 273 = 194K

= 44.01 × 29.86 × 194

= 254942.89J/1000

= 254.943kJ/mol